# Azumaya algebras and Barr’s Theorem

### v2

@@
Thierry Coquand Computer Science and Engineering Department, University of Gothenburg, Sweden, coquand@chalmers.se Henri Lombardi Université Marie et Louis Pasteur, CNRS, LmB, FR-25000 Besançon, France, henri.lombardi@univ-fcomte.fr, stefan.neuwirth@univ-fcomte.fr Stefan Neuwirth Université Marie et Louis Pasteur, CNRS, LmB, FR-25000 Besançon, France, henri.lombardi@univ-fcomte.fr, stefan.neuwirth@univ-fcomte.fr
(May 17, 2025)
@@ Introduction
A fundamental invariant associated to any commutative field $F$ is its Brauer group, whose elements are division algebras over $F$, up to isomorphism. Alternatively an element of the Brauer group can be described as a central simple algebra, up to Morita equivalence, and the group operation is then the tensor product. Two important characterisations of the notion of central simple algebra are the following:
@@
    an algebra $A$ which becomes a matrix algebra over an algebraic (resp. separable) extension of $F$.
@@ 1.1 Theory of local rings
then we get the theory of local rings that are residually discrete, i.e. where the property of being invertible is now decidable, and the predicate $\mathsf{m}(x)$ corresponds to the unique maximal ideal of the ring. In this theory, we can express in a coherent way that the ring is Henselian by stating that for any monic polynomial $P$, if we have $u$ such that $P(u)\in\mathsf{m}$ and $P^{\prime}(u)\notin\mathsf{m}$, we can find $\alpha$ such that $P(\alpha)=0$ and $\alpha-u\in\mathsf{m}$. Equivalently [ALP], any monic polynomial of degree $l+1$ which is of the form $X^{l}(X-1)$ mod. $\mathsf{m}$ has a root which is equal to $1$ mod. $\mathsf{m}$. The work [ALP] shows in an intuitionistic setting for this theory the lifting property for idempotents of a finitely generated algebra over a Henselian local ring.
@@ 1.2 Theory of separably closed local rings
If $R$ is a ring, we define the Zariski spectrum of $R$ to be the distributive lattice generated by symbols $D(a)$ and relations
@@
We write $D(a_{1},\dots,a_{n})$ for $D(a_{1})\vee\dots\vee D(a_{n})$.
One way to build this lattice is to define $D(a_{1},\dots,a_{n})$ to be the radical of the ideal $(a_{1},\dots,a_{n})$.
@@
It is also direct to see that $D(a,b)=D(a+b,ab)$ and more generally that
@@
For the formulation of this theory of strictly Henselian rings, we need the notion of universal decomposition algebra $L=R[x_{1},\dots,x_{n}]=R[X_{1},\dots,X_{n}]/I$ of a monic polynomial
@@
in $R[X]$, where $I$ is the ideal generated by $u_{1}-\sigma_{1},~{}\dots,~{}u_{n}-\sigma_{n}$, where $\sigma_{1},\dots,\sigma_{n}$ in $R[X_{1},\dots,X_{n}]$ are the elementary symmetric polynomials. We know that $L=R[x_{1},\dots,x_{n}]$ is freely generated as a module over $R$ and it follows from this that $L$ is a faithfully flat extension of $R$.
@@
Over a residually discrete local ring with residue field $F$, to be unramifiable means that the polynomial $f$ has a simple root in some extension of $F$.
@@
Let $R$ be a local ring residually discrete of maximal ideal $\mathsf{m}$ with residue field $F=R/\mathsf{m}$. If $f$ is a monic polynomial in $R[X]$ which has a simple root in some nontrivial $F$-algebra $A$, then $f$ is unramifiable.
@@
We have in $R[x_{1},\dots,x_{n}]$, the decomposition algebra of $g$,
@@
in $R[x_{1},\dots,x_{n}]$. Since the decomposition algebra is faithfully flat, we get
@@
Let $P$ be a monic polynomial which has residually a simple root $u$. We show by induction on the degree of $P$ that $P$ has a root in $R$ which is equal to $u$ mod. $\mathsf{m}$. The polynomial $P$ is unramifiable by Lemma 1.1, since it has residually a simple root, and it has a root $a$. If $a\neq u$ mod. $\mathsf{m}$ then we write $P=(X-a)Q$, and $u$ is residually a simple root of $Q$. ∎
@@ 2 Some lemmas about Henselian rings
Let $A=R[x_{1},\dots,x_{n}]$ be the universal decomposition algebra of $P$. This is a free $R$-module over $R$ of dimension $n!$ and the symmetric group acts on it [LQ, Fact 3-4.4]. We may identify $R[x]$ with the subalgebra $R[x_{1}]$ and we have $e(x_{1})$ in $R[x_{1}]$ which is idempotent mod. $\mathsf{m}R[x_{1}]$. We show that we have $u_{1}$ in $R[x_{1}]$ which is idempotent in $R[x_{1}]$ and equal to $e(x_{1})$ mod. $\mathsf{m}R[x_{1}]$.
The product $e(x_{1})\cdots e(x_{n})$ is in $R$ and it is an idempotent mod. $\mathsf{m}$, so it is $1$ or 00 mod. $\mathsf{m}$. If it is $1$, we have $e(x_{1})=1$ mod. $\mathsf{m}R[x_{1}]$ and we can take $u_{1}=1$.
@@
Let $u$ be a vector column of generators of $M$, then we can find a vector $v$ in $N$ and a matrix $P$ with elements in $\mathsf{m}$ such that $u=v+Pu$. We then have $(I-P)u=v$ and hence the elements of $v$ are generators of $M$ since $\det(I-P)$ is invertible. ∎
@@
Let $A$ be an $R$-algebra which is finitely generated as an $R$-module, and $u$ in $A$ which is invertible mod. $\mathsf{m}A$, then $u$ is invertible in $A$.
@@
We let $R[x]=R[X]/(P)$. We can find an idempotent $e$ in $R[x]$ mod. $\mathsf{m}R[x]$ such that $(f)=(e),~{}(g)=(1-e)$ mod. $\mathsf{m}R[x]$ [ALP, Prop. 3.11]. Using Lemma 2.1, we may assume $e$ idempotent in $R[x]$. Let $n$ (resp. $m$) be the degree of $f$ (resp. $g$). Then $e,ex,\dots,ex^{m-1}$ and $(1-e),(1-e)x,\dots,(1-e)x^{n-1}$ generate $R[x]$ as an $R$-module mod. $\mathsf{m}$. By Lemma 2.2, they generate the $R$-module $R[x]$. In particular, $ex^{m}$ belongs to the $R$-module generated by $e,ex,\dots,ex^{m-1}$ (since $e$ is idempotent in $R[x]$) and we find in this way $G$ monic of degree $m$ such that $eG(x)=0$ in $R[x]$. Similarly we find $F$ monic of degree $n$ such that $(1-e)F(x)=0$ in $R[x]$. We then have $F(x)G(x)=F(x)eG(x)=0$ and hence $P$ divides $FG$. But they have the same degree, hence $P=FG$. ∎
@@
We consider the subalgebra $R[a]$ of $A$. Since $A$ is finitely generated as an $R$-module we first find $F$ monic in $R[X]$ such that $F(a)=0$. We decompose $F=X^{n}(1-X)^{m}h$ mod. $\mathsf{m}R[X]$, with $h$ monic and $h(0),~{}h(1)$ not in $\mathsf{m}$. If $n=0$ we have $F(0)=h(0)$ mod. $\mathsf{m}$. Since $F(a)=0$, we get that $a$ is invertible mod. $\mathsf{m}A$, and hence invertible in $A$ using Corollary 2.1, and we can take $e=1$. Similarly, if $m=0$, we get that $1-a$ is invertible and we can take $e=0$.
In the remaining case, we can find using Lemma 2.3 a decomposition in monic polynomials $F=PQH$ in $R[X]$ with $P=X^{n},~{}Q=(1-X)^{m},~{}H=h$ mod. $\mathsf{m}$. We also have $U$ and $V$ in $R[X]$ such that $X(1-X)U+HV=Res(X(1-X),H)=1$ [LQ, Fact 3-7.1], [ALP, Lemma 3.10], and so $H(a)V(a)=1$ mod. $\mathsf{m}A$, and so $H(a)$ is invertible mod. $\mathsf{m}A$ and so invertible in $A$ using Corollary 2.1. Since $F(a)=P(a)Q(a)H(a)=0$, we have $P(a)Q(a)=0$ in $A$. We also have $P(a)=a$ and $Q(a)=1-a$ mod. $\mathsf{m}A$ and so $P(a)+Q(a)=1$ mod. $\mathsf{m}A$ and is also invertible in $A$, by Corollary 2.1, with inverse $\mu$. We can then take $e=\mu P(a)$. ∎
@@ 3 Azumaya algebras
    $A$ is projective of finite type;
@@
Note that we allow the trivial algebra to be an Azumaya algebra.
@@
$A$ is an Azumaya algebra over $k$ if, and only if, we can build a finite tree of root $k$ using the following, where all leaves are matrix algebras:
@@
If we have such a tree starting from $k$, notice that each extension $S\rightarrow S[X]/(P)$ for $P$ monic is faithfully flat, and $S\rightarrow\prod_{i}S[1/u_{i}]$ if $1=(u_{1},\dots,u_{n})$ is also faithfully flat. It follows then that $A$ is finite projective [LQ, Thm. VIII-6.7]. The map $A\otimes_{k}A^{op}\rightarrow End_{k}(A)$ is then a bijection since it becomes a bijection by faithfully flat extension.
The converse follows from the following two results.
@@
We simplify the argument in Milne’s course notes ([Milne, Chapter IV, Proposition 1.6]), using the constructive development in [ALP] and [CLN]. If $\mathsf{m}$ is the maximal ideal of $R$, we have that $A$ is split over $R/\mathsf{m}$ ([CLN]), and hence we have a matrix algebra decomposition $e_{ij}$ mod. $\mathsf{m}A$. We can then lift the idempotent $e_{11}$ mod. $\mathsf{m}A$ to some idempotent $e$ in $R$ using Theorem 2.1. We have that $Ae$ is projective and hence free since $R$ is a local ring. The map $A\rightarrow End_{R}(Ae)$ is then an isomorphism, since it is a map between two free $R$-modules of the same dimension, and it is residually an isomorphism. ∎
@@
For a given Azumaya algebra, the conclusion is coherent, and the hypothesis that $A$ is an Azumaya algebra is formulated in a coherent theory. Since it is proved with the extra hypothesis that $R$ is residually discrete, the syntactic version of Barr’s Theorem [^3] We can actually follow the argument in [Joh] and apply Friedman-Dragalin’s translation [Avi] w.r.t. to the proposition expressing that $A$ is a matrix algebra. (constructively valid) shows that it can be proved without this hypothesis. ∎
Theorem 3.1 follows by interpreting this result in the classifying topos of the theory of strictly Henselian $k$-algebra.
@@
If $R$ is a local ring and $A$ a matrix algebra of rank $>0$ [LQ, Def. X-2.2] then any automorphism of $A$ is of the form $x\mapsto axa^{-1}$ for some $a$ in $GL_{n}(A)$.
@@
If $A$ is an Azumaya algebra, of rank $>0$ as a projective module over a ring $k$, and $\psi$ is an automorphism of $A$, then $M=\{a\in A~{}|~{}\forall_{x}ax=\psi(x)a\}$ is a projective $k$-module of rank $1$.
@@
We use Theorem 3.1. The statement holds on the leaves ($M$ is even free of rank $1$ on the leaves by Lemma 3.2) and descends to the root by tree induction using the local-global principle and the descent principle w.r.t. faithfully flat extension for projective modules ([LQ, Thm. VIII-6.7]). ∎
@@
Claude Quitté suggested the following refinement.
@@
If $A$ is an Azumaya algebra, of rank $>0$ as a projective module over a ring $k$, and $\psi$ is an automorphism of $A$, then
@@
We have on the leaves $MN\subseteq k1$ and this statement descends to the root by tree induction. This can be used to define a map $M\rightarrow Hom(N,k),~{}x\mapsto(y\mapsto xy)$. This map is an isomorphism on the leaves, and this also descends to the root by tree induction. Finally, $M$ and $N$ are direct factors on the leaves, and this fact descends as well. ∎

### v3

@@
Thierry Coquand Henri Lombardi Stefan Neuwirth
(November 24, 2025)
@@ Introduction
A fundamental invariant associated to any commutative field $F$ is its Brauer group, whose elements are division algebras over $F$, up to isomorphism. Alternatively an element of the Brauer group can be described as a central simple algebra, up to Morita equivalence, and the group operation is then the tensor product. Two important characterisations of the notion of central simple algebra are the following [^1] We study these characterisations in a constructive setting in the reference [CLN].:
@@
    an algebra $A$ which becomes a matrix algebra over an algebraic extension of $F$, or, equivalently, which becomes a matrix algebra over an algebraic and separable extension of $F$.
@@ 1.1 Theory of local rings
then we get the theory of local rings that are residually discrete [^2] The terminology “discrete” for a set with a decidable equality, while conflicting with some topological intuitions, is well established in constructive mathematics since the work of Bishop [Bish], and we follow this usage in this paper., i.e. where the property of being invertible is now decidable, and the predicate $\mathsf{m}(x)$ corresponds to the unique maximal ideal of the ring. In this theory, we can express in a coherent way that the ring is Henselian by stating that for any monic polynomial $P$, if we have $u$ such that $P(u)\in\mathsf{m}$ and $P^{\prime}(u)\notin\mathsf{m}$, we can find $\alpha$ such that $P(\alpha)=0$ and $\alpha-u\in\mathsf{m}$. Equivalently [ALP], any monic polynomial of degree $l+1$ which is of the form $X^{l}(X-1)$ mod. $\mathsf{m}$ has a root which is equal to $1$ mod. $\mathsf{m}$. The work [ALP] shows in an intuitionistic setting for this theory the lifting property for idempotents of a finitely generated algebra over a Henselian local ring.
@@ 1.2 Theory of separably closed local rings
If $R$ is a ring, and $a_{1},\dots,a_{n}$ elements of $R$, we write $(a_{1},\dots,a_{n})$ for the ideal of $R$ generated by $a_{1},\dots,a_{n}$.
We define, following Joyal [Joy], the Zariski spectrum of $R$ as the universal support $D:R\rightarrow Z(R)$ over $R$. A support is a map $d:R\rightarrow M$ into a bounded distributive lattice $M$ satisfying the relations
@@
We write $d(a_{1},\dots,a_{n})$ for $d(a_{1})\vee\dots\vee d(a_{n})$. It is direct that $d(a)\leqslant d(a_{1},\dots,a_{n})$ as soon as $a$ belongs to the radical of the ideal $(a_{1},\dots,a_{n})$. Conversely, it is elementary to check that we get a support if we define $D(a_{1},\dots,a_{n})$ to be the radical of the ideal $(a_{1},\dots,a_{n})$, and this is a concrete way to build the universal support.
@@
It is also direct, for any support $d$, to see that $d(a,b)=d(a+b,ab)$ and more generally that
@@
For the formulation of this theory of strictly Henselian rings, we need the notion of universal splitting algebra [LQ, III, 4] $L=R[x_{1},\dots,x_{n}]=R[X_{1},\dots,X_{n}]/I$ of a monic polynomial
@@
in $R[X]$, where $I$ is the ideal generated by $u_{1}-\sigma_{1},~\dots,~u_{n}-\sigma_{n}$, where $\sigma_{1},\dots,\sigma_{n}$ in $R[X_{1},\dots,X_{n}]$ are the elementary symmetric polynomials. We know, using Cauchy modules [LQ, III, 4.3], that $L=R[x_{1},\dots,x_{n}]$ is freely generated of rank $n!$ as a module over $R$ and it follows from this that $L$ is a faithfully flat extension of $R$.
For the notion of faithfully flat extension, we follow the definition in [LQ, VIII, 6]: it is an $R$-algebra $i:R\rightarrow S$ such that
  • if $u$ is a row vector in $A^{n}$ and $v$ a column vector in $B^{n}$ such that $i(u)v=0$ then we can find a matrix $M$ with coefficients in $A$ and $w$ such that $uA=0$ and $v=i(A)w$;
  • if $i(a)$ belongs to the ideal $(i(a_{1}),\dots,i(a_{n}))$ in $B$ then $a$ belongs to the ideal $(a_{1},\dots,a_{n})$ in $A$.
It then follows [LQ, VIII, 6] that $i$ is injective. It is also clear from this definition that $S$ is faithfully flat over $R$ when it is a finite free $R$-module of rank >0>0.
@@
Over a residually discrete local ring with residue field $F$, to be unramifiable means that the polynomial $f$ has a simple root in some extension of $F$. This holds in classical mathematics, but also constructively with a dynamical interpretation of extensions of $F$.
@@
Let $R$ be a local ring, residually discrete with maximal ideal $\mathsf{m}$ and residue field $F=R/\mathsf{m}$. If $f$ is a monic polynomial in $R[X]$ which has a simple root in some nontrivial $F$-algebra AA, then $f$ is unramifiable.
@@
We have in $R[x_{1},\dots,x_{n}]$, the universal splitting algebra of $g$,
@@
in $R[x_{1},\dots,x_{n}]$. Since the universal splitting algebra is faithfully flat, we get
@@
Let $P$ be a monic polynomial which has residually a simple root $u$. We show by induction on the degree of $P$ that $P$ has a root in $R$ which is equal to $u$ mod. $\mathsf{m}$. The polynomial $P$ is unramifiable by Lemma 1.1, since it has residually a simple root, and it has a root $a$. If $a\neq u$ mod. $\mathsf{m}$ then we write P=(Xa)QP=(X-a)Q, and $u$ is residually a simple root of $Q$. ∎
@@ 2 Some lemmas about Henselian rings
Let $A=R[x_{1},\dots,x_{n}]$ be the universal splitting algebra of $P$. This is a free $R$-module over $R$ of dimension $n!$ and the symmetric group acts on it [LQ, Fact III-4.4]. We may identify $R[x]$ with the subalgebra $R[x_{1}]$ and we have $e(x_{1})$ in $R[x_{1}]$ which is idempotent mod. $\mathsf{m}R[x_{1}]$. We show that we have $u_{1}$ in $R[x_{1}]$ which is idempotent in $R[x_{1}]$ and equal to $e(x_{1})$ mod. $\mathsf{m}R[x_{1}]$.
The product $e(x_{1})\cdots e(x_{n})$ is in $R$ and it is an idempotent mod. $\mathsf{m}$, so it is $1$ or 0 mod. $\mathsf{m}$. If it is $1$, we have $e(x_{1})=1$ mod. $\mathsf{m}R[x_{1}]$ and we can take $u_{1}=1$.
@@
Let $u$ be a vector column of generators of $M$: then we can find a vector $v$ in $N$ and a matrix $P$ with elements in $\mathsf{m}$ such that $u=v+Pu$. We then have (IP)u=v(I-P)u=v and hence the elements of $v$ are generators of $M$ since $\det(I-P)$ is invertible. ∎
@@
Let $A$ be an $R$-algebra which is finitely generated as an $R$-module, and $u$ in $A$ which is invertible mod. $\mathsf{m}A$. Then $u$ is invertible in $A$.
@@
We let $R[x]=R[X]/(P)$. We can find an idempotent $e$ in $R[x]$ mod. $\mathsf{m}R[x]$ such that (f)=(e),(g)=(1e)(f)=(e),~(g)=(1-e) mod. $\mathsf{m}R[x]$ [ALP, Prop. 3.11]. Using Lemma 2.1, we may assume $e$ idempotent in $R[x]$. Let $n$ (resp. $m$) be the degree of $f$ (resp. $g$). Then $e,ex,\dots,ex^{m-1}$ and $(1-e),(1-e)x,\dots,(1-e)x^{n-1}$ generate $R[x]$ as an $R$-module mod. $\mathsf{m}$. By Lemma 2.2, they generate the $R$-module $R[x]$. In particular, $ex^{m}$ belongs to the $R$-module generated by $e,ex,\dots,ex^{m-1}$ (since $e$ is idempotent in $R[x]$) and we find in this way $G$ monic of degree $m$ such that $eG(x)=0$ in $R[x]$. Similarly we find $F$ monic of degree $n$ such that (1e)F(x)=0(1-e)F(x)=0 in $R[x]$. We then have $F(x)G(x)=F(x)eG(x)=0$ and hence $P$ divides $FG$. But they have the same degree, hence $P=FG$. ∎
@@
We consider the subalgebra $R[a]$ of $A$. Since $A$ is finitely generated as an $R$-module we first find $F$ monic in $R[X]$ such that $F(a)=0$. We decompose $F=X^{n}(1-X)^{m}h$ mod. $\mathsf{m}R[X]$, with $h$ monic and $h(0)$, $h(1)$ not in $\mathsf{m}$. If $n=0$ we have $F(0)=h(0)$ mod. $\mathsf{m}$. Since $F(a)=0$, we get that $a$ is invertible mod. $\mathsf{m}A$, and hence invertible in $A$ using Corollary 2.1, and we can take $e=1$. Similarly, if $m=0$, we get that 1a1-a is invertible and we can take $e=0$.
In the remaining case, we can find using Lemma 2.3 a decomposition in monic polynomials $F=PQH$ in $R[X]$ with $P=X^{n},~Q=(1-X)^{m},~H=h$ mod. $\mathsf{m}$. We also have $U$ and $V$ in $R[X]$ such that X(1X)U+HV=Res(X(1X),H)=1X(1-X)U+HV=Res(X(1-X),H)=1 [LQ, Fact 3-7.1], [ALP, Lemma 3.10], and so $H(a)V(a)=1$ mod. $\mathsf{m}A$, and so $H(a)$ is invertible mod. $\mathsf{m}A$ and so invertible in $A$ using Corollary 2.1. Since $F(a)=P(a)Q(a)H(a)=0$, we have $P(a)Q(a)=0$ in $A$. We also have $P(a)=a$ and Q(a)=1aQ(a)=1-a mod. $\mathsf{m}A$ and so $P(a)+Q(a)=1$ mod. $\mathsf{m}A$ and is also invertible in $A$, by Corollary 2.1, with inverse $\mu$. We can then take $e=\mu P(a)$. ∎
@@ 3 Azumaya algebras
    AA is projective of finite type;
@@
Note that we allow the trivial algebra to be an Azumaya algebra. Note also that any matrix algebra is Azumaya.
One can equivalently describe the algebra structure by a bilinear map $\mu:A\times A\rightarrow A$; its scalar extension along $k\rightarrow S$ is the unique $S$-algebra $A_{S}$ through which every bilinear map extending $\mu$ factors. Thus all subsequent arguments can be phrased without explicit tensor notation.
Over a finite free $k$-module, to give an algebra structure is represented concretely by a multiplication table. A module $A$, projective of finite type over $k$, is given concretely by a projection matrix over $k$ [LQ], and we can similarly represent the algebra structure by a multiplication table (so $m^{3}$ elements of $k$ if the projection matrix has finite size $m$). We can then compute [LQ, X, Theorem 1.6] comaximal elements $s_{1},\dots,s_{l}$ such that $A_{s_{i}}$ is free over $k_{s_{i}}$, and the second condition for being an Azumaya algebra becomes then that some determinants $\delta_{1},\dots,\delta_{l}$ are invertible [^5] After localisation to $s_{j}$ the Azumaya algebra becomes free of rank $r_{i}=n_{i}^{2}\leqslant m$.. From this discussion follows that if $A$ is an Azumaya algebra, given by a multiplication table, and we have a $k$-algebra $k\rightarrow S$ then the algebra $A_{S}$ over $S$, with the same multiplication in $S$, is also an Azumaya algebra.
@@
AA is an Azumaya algebra over $k$ if, and only if, we can build a finite tree with root $k$ using the following, where at all leaves all scalar extensions $A_{S}$ are matrix algebras:
@@
We prove first that if we have such a tree then the algebra $A$ is Azumaya. If we have such a tree starting from $k$, notice that each extension $S\rightarrow S[X]/(P)$ for $P$ monic is faithfully flat, and $S\rightarrow\prod_{i}S[1/u_{i}]$ if $1=(u_{1},\dots,u_{n})$ is also faithfully flat. It follows then that $A$ is finite projective [LQ, Thm. VIII-6.7]. The map $A\otimes_{k}A^{op}\rightarrow End_{k}(A)$ is then a bijection since it becomes a bijection by a faithfully flat extension. (As we saw before, to be a bijection is witnessed by the fact that some determinants are invertible, and an element is invertible if it becomes invertible by a faithfully flat extension.)
For the converse, we look at the algebra $A$ (given concretely by its multiplication table) inside the étale topos over $k$, using the site defined in the first section. (This is similar to the constructive use of the algebraic closure in [MC].) The generic ring $R$, defined by $R(S)=S$ for $S$ a finitely presented $k$-algebra, is, in the internal logic of this topos, a local strictly Henselian ring. The algebra $A$ defines then an $R$-algebra $A_{R}(S)=A_{S}$, which is an Azumaya algebra, given by the same multiplication table as the one of $A$. If we can prove in intuitionistic logic that an Azumaya algebra over a strictly Henselian ring is a matrix algebra, we will get, using the interpretation of intuitionistic logic in a sheaf model [Joh, MC], a covering tree of $k$ which gives Theorem 3.1.
Internally in the classifying topos of stricly Henselian $k$-algebras, the object $R$ is local strictly Henselian and the claim that $A_{R}$ is a matrix algebra is a coherent statement, provable intuitionistically. Such a proof yields externally a finite covering tree of $k$ satisfying the conditions above.
The converse is a consequence of the following two results.
@@
We simplify the argument in Milne’s course notes ([Milne, Chapter IV, Proposition 1.6]), using the constructive development in [ALP] and [CLN]. If $\mathsf{m}$ is the maximal ideal of $R$, we have that $A$ is split over $R/\mathsf{m}$ ([CLN]), and hence we have a matrix algebra decomposition $e_{ij}$ mod. $\mathsf{m}A$. We can then lift the idempotent $e_{11}$ mod. $\mathsf{m}A$ to some idempotent $e$ in $R$ using Theorem 2.1. We have that $Ae$ is projective and hence free since $R$ is a local ring. The map $A\rightarrow End_{R}(Ae)$ is then an isomorphism, since it is a map between two free $R$-modules of same dimension, and it is residually an isomorphism. ∎
The next proof uses Barr’s Theorem, similar to the way it is used in [Penon, Wraith], but the difference is that we only need a constructively valid version of this theorem.
@@
For a given Azumaya algebra, the conclusion is coherent [Joh, MC], and the hypothesis that $A$ is an Azumaya algebra is formulated in a coherent theory. Since it is proved with the extra hypothesis that $R$ is residually discrete, the syntactic version of Barr’s Theorem [^6] We can actually follow the argument in [Joh] and apply Friedman-Dragalin’s translation [Avi] with respect to the proposition expressing that $A$ is a matrix algebra. (constructively valid) shows that it can be proved without this hypothesis. ∎
Theorem 3.1 follows by interpreting this result in the classifying topos of the theory of strictly Henselian $k$-algebras.
@@
If $R$ is a local ring and $A$ a matrix algebra of rank $\geqslant 1$ [LQ, Def. X-2.2] then any automorphism of $A$ is of the form $x\mapsto axa^{-1}$ for some $a$ in $GL_{n}(A)$.
@@
For the next result, we use the local-global principle and the descent principle with respect to faithfully flat extension for projective modules ([LQ, Theorem VIII-6.7]). For completeness, we recall the statement. Note that the proof in [LQ] is constructive, using the concrete definition of faithfully flat extension and the fact that, also constructively, a module is finite projective if, and only if, it is finitely presented and flat.
Theorem 3.2 (Descent).
Let $A\rightarrow B$ be a faithfully flat extension and $M$ an $A$-module. If $M\otimes_{A}B$ is finite projective, then so is $M$. If we have $1=(s_{1},\dots,s_{n})$ in $A$ and each $M_{s_{i}}$ is finite projective over $A_{s_{i}}$ then so is $M$.
@@
If $A$ is an Azumaya algebra, of rank $\geqslant 1$ as a projective module over a ring $k$, and $\psi$ is an automorphism of $A$, then $M=\{a\in A~|~\forall_{x}\,ax=\psi(x)a\}$ is a projective $k$-module of rank $1$.
@@
We use Theorem 3.1. The statement holds on the leaves ($M$ is even free of rank 11 on the leaves by Lemma 3.2) and descends to the root by tree induction using Theorem 3.2. ∎
@@
Claude Quitté has suggested the following refinement.
@@
If $A$ is an Azumaya algebra, of rank $\geqslant 1$ as a projective module over a ring $k$, and $\psi$ is an automorphism of $A$, then
@@
We have on the leaves $M_{S}N_{S}\subseteq S1$ and this statement descends to the root by tree induction. This can be used to define a map $M\rightarrow Hom(N,k),~x\mapsto(y\mapsto xy)$. This map is an isomorphism on the leaves, and this also descends to the root by tree induction. Finally, $M$ and $N$ are direct factors on the leaves, and this fact descends as well. ∎